If PQ = x – 5, QR = 2x – 1, and PR = 21, find x. x =


1
Expert's answer
2021-09-01T23:47:32-0400

Let "PQ = x - 5,\ QR = 2x - 1," and "PR = 21." Taking into account that "PR=PQ+QR," we get that "x-5+2x-1=21," and hence "3x=27." It follows that "x=9."